OpenStudy (anonymous):

What is the simplified square root of 63? showing all your work. No decimals allowed

OpenStudy (anonymous):

3√7 maybe

OpenStudy (anonymous):

√63 = √(3^2 * 7) = √(3^2) * √7 = 3 * √7 = 3√7

OpenStudy (anonymous):

thanks